OpenStudy (anonymous):

integrate 3/((x+1)^2)

OpenStudy (anonymous):

\[\int\limits_{}^{}(3/(x+1)^{2}))dx=-3/(x+1)+c\] it's very easy ........

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

(he used a simple u substitution)